OpenStudy (anonymous):

Am I right? Take a quick look? You purchase a house in 1995 for $95000. The house increases per year you own it. Write the exponential model for the value of the house. A= 95000(1.03)^t Use the model to find the value of the house after 10 years. A= 127,672.056 How long will it take until the value of the house reaches $50,000? T= 1.533 years

OpenStudy (lyrae):

If the value of the house increases by 3% per year, 1 and 2 look correct. For the third one, how can the house reach a value of $50,000 if it's alrady worth $95,000? Are you sure they don't actually mean $50,000 value increase?

OpenStudy (anonymous):

Dude i wish i knew but thats what the question says. Not much i can do about that but yeah they probably meant that.

OpenStudy (lyrae):

Then, how did you get T= 1.533 years?

OpenStudy (anonymous):



OpenStudy (anonymous):

Thats the equation and got that.answer

OpenStudy (lyrae):

\[50000=95000 \times (1.03)^t \]\[\frac{ 50000 }{ 95000} = \times (1.03)^t\]\[\large t = \frac{ \lg(\frac{ 50000 }{ 95000}) }{ \lg(1.03) } \approx -21.71\]

OpenStudy (lyrae):

Disregard that multiplication sign on the right side in the middle one

OpenStudy (lyrae):

Anyway I think a $50000 will produce a better answer because answers in negative time usually's not correct.

OpenStudy (anonymous):

Actually I recalculated it and i got a negative this time so yeah. . .. .

OpenStudy (lyrae):

a $50000 value increase*

OpenStudy (anonymous):

So would it be like no solution or something.

OpenStudy (lyrae):

Yeah, depending on how you interpret the question. Either that, or if it's a $50000 value incerase then you'll get the equation \[95000+50000= 95000 \times 1.03^t \]

OpenStudy (lyrae):

which has a positive soulution

OpenStudy (anonymous):

Alright. Thanks.

OpenStudy (lyrae):

Yw, no problem :)